Experiment 10 Prelaboratory Assignment Vinegar Ana
Experiment 10 Prelaboratory Assignment Vinegar AnaNeed help with this prelab please
Experiment 10 Prelaboratory Assignment Vinegar Ana
Answer #1

1) We have 5% acetic acid solution, i.e, 100 gm vinegar solution contains 5 gm acetic acid.

Therefore, 1000 gm vinegar solution will have (5*1000)/100 gm = 50 gm acetic acid.

Now, density of the solution is 1 gm/mL; therefore, volume of 1000 gm solution is (1000 gm)/(1 gm/mL) = 1000 Ml = 1 L.

Thus 1 L vinegar solution contains 50 gm acetic acid.

Molar mass of acetic acid = 60.05 gm/mol.

Therefore, moles of acetic acid in 50 gm sample = (50 gm)*(1 mole/60.05 gm) = 0.8326 mole.

Molarity of the supplied vinegar/acetic acid solution = 0.8326 mole/1 L = 0.8326 M (since the volume of the solution is 1 L).

The reaction between acetic acid and NaOH is

CH3COOH + NaOH ---------> CH3COO-Na+ + H2O

The reaction is a 1:1 reaction. Therefore, using the law of molarity, we have

(25 mL)*(0.10 M) = V*(0.8326 M)

===> V = (25*0.10)/0.8326 = 3.002 ≈ 3.0

The volume of acetic acid required = 3.0 mL (ans)

2a) It is easy to identify a white mark or reading against a black background rather than reading a white mark against a white background. Hence, a white card with a black mark is used to make the reading easier and more prominent.

b) The titre (the solution in the burette) can stick to the walls of the container in which the titration is being carried out and thus the wait is ideal to make sure that the titrant droplets have fallen into the solution and are not sticking to the walls. Such sticking to the walls will give an erroneous result.

c) A temporary color change occurs possibly due to a local reaction between the titre (solution in the burette) and the titrant (solution in the beaker) to afford a complex. Most titration reactions are pH dependent; hence if the correct pH is not attained, the said complex will decompose. At the equivalence point, the solution has the correct pH to stabilize the said complex. The complex will no longer decompose at the end point and the color change will persist. Thus, the 30 second wait is required to ascertain the correct equivalence point.

3) The citric acid concentration is 0.5% by mass, i.e, 100 gm lemon juice contains 0.5 gm citric acid.

Therefore, 3.71 gm lemon juice contains (3.71 gm)*(0.5 gm/100 gm) = 0.01855 gm citric acid.

The density of lemon juice is 1gm/mL; therefore, volume of 3.71 gm lemon juice = 3.71 gm/(1.0 gm/mL) = 3.71 mL.

So, 3.71 mL lemon juice contains 0.01855 gm citric acid.

Molar mass of citric acid = 190.12 gm/mol. Therefore, moles of citric acid in 0.01855 gm citric acid = (0.01855 gm)*( 1 mole/190.12 gm) = 9.757*10-5 mole.

Now, the reaction between NaOH and citric acid is HA + NaOH ------> NaA + H2O, i.e, 1:1 ratio. Hence, 9.757*10-5 mole citric acid requires 9.757*10-5 mole NaOH for neutralization.

The molarity of the given NaOH solution is 0.0100 M (= 0.0100 mole/L).Let the volume required be V mL.

Thus,

V*(1 L/1000 mL)*(0.0100 mole/L) = 9.757*10-5 mole

===> V = (9.757*10-5*1000)/0.0100 mL = 9.757 mL ≈ 9.8 mL

The volume of NaOH required = 9.8 mL (ans)

5a) (B) (3) Volume of NaOH used = (final reading – initial reading) = (25.40 – 3.70) mL = 21.70 mL.

(5) Moles of NaOH added = (volume in L)*(concentration in moles/L) = 21.70 mL*(1 L/1000 mL)*(0.0940 mol/L) = 2.0398*10-3 mole ≈ 2.04*10-3 mol

(6) Moles of acetic acid in vinegar = 2.04*10-3 mol (since the reaction between acetic acid and NaOH is a 1:1 reaction).

(7) Mass of CH3COOH in vinegar = (moles of CH3COOH)*(molar mass of CH3COOH) = (2.04*10-3 mol)*(60.05 gm/1 mol) = 0.1224 gm

(8) Percent by mass of CH3COOH in vinegar = (0.1224 gm)/(3.06 gm)*100 = 4.00 ≈

5b) The three values are: 4.00%, 5.01% and 4.66%. The average value is (4.00 + 5.01 + 4.66)/3 % = 4.556% ≈ 4.56%